# A Bayesian Approach to Segmentation with Noisy Labels via Spatially Correlated Distributions

### v1

@@ Abstract
In semantic segmentation, the accuracy of models heavily depends on the high-quality annotations. However, in many practical scenarios such as medical imaging and remote sensing, obtaining true annotations is not straightforward and usually requires significant human labor. Relying on human labor often introduces annotation errors, including mislabeling, omissions, and inconsistency between annotators. In the case of remote sensing, differences in procurement time can lead to misaligned ground truth annotations. These label errors are not independently distributed, and instead usually appear in spatially connected regions where adjacent pixels are more likely to share the same errors. To address these issues, we propose an approximate Bayesian estimation based on a probabilistic model that assumes training data includes label errors, incorporating the tendency for these errors to occur with spatial correlations between adjacent pixels. Bayesian inference requires computing the posterior distribution of label errors, which becomes intractable when spatial correlations are present. We represent the correlation of label errors between adjacent pixels through a Gaussian distribution whose covariance is structured by a Kac-Murdock-Szegö (KMS) matrix, solving the computational challenges. Through experiments on multiple segmentation tasks, we confirm that leveraging the spatial correlation of label errors significantly improves performance. Notably, in specific tasks such as lung segmentation, the proposed method achieves performance comparable to training with clean labels under moderate noise levels. Code is available at https://github.com/pfnet-research/Bayesian_SpatialCorr.
Keywords:
Segmentation Noisy label Noisy annotation Bayesian inference Spatial correlation.
@@ 1 Introduction
Semantic segmentation, which involves classifying each pixel in an image into one of several classes, is a crucial task in computer vision. In supervised learning, the accuracy of segmentation models critically depends on the quality of the annotations in the training data. However, obtaining truly accurate pixel-level annotations is challenging in many practical applications. Even when expert annotators are employed, errors, omissions, and subjectivity in interpretation are inevitable, leading to inconsistencies in datasets. In particular, high inter- and intra-annotator variability is widely reported in medical imaging, where experts may have differing interpretations of the same structures. For instance, [41, 19, 27, 38, 27] highlight significant discrepancies among expert annotators; some delineate structures more generously, while others prefer more conservative annotations. The observer-dependent annotations exacerbate label noise in supervised learning. Label noise is also a critical issue in remote sensing, where determining ground truth labels often requires field surveys over large and sometimes inaccessible regions [7, 26, 6]. Due to the logistical and economic challenges of large-scale ground truth collection, researchers frequently rely on automatic labeling systems, which may introduce systematic errors. Additionally, high-quality annotated datasets remain a critical bottleneck for supervised learning, particularly in remote sensing applications where annotations are often repurposed across different types of satellite images. For example, the OpenEarthMap dataset [35] was created by manually annotating high-resolution optical satellite images for semantic segmentation. However, these annotations are sometimes reused for synthetic aperture radar (SAR) imagery, despite differences in resolution and capture conditions [12, 42, 23]. Additionally, changes in artificial structures or variations in land cover further contribute to label inconsistencies [8, 12].
@@
Various approaches have been proposed to mitigate the adverse effects of noisy labels. Some methods attempt to stop training early to prevent the network from overfitting to noise and generating unreliable pseudo-labels [21, 22], while others modify the loss function to be more robust against large errors [9]. Although these techniques can reduce the influence of noisy annotations, they do not fundamentally address the core reason why the standard supervised learning framework fails in the presence of noisy labels.
@@
We propose a method that directly tackles the root cause of this issue. In supervised learning for segmentation models, training typically reduces to optimizing the cross-entropy loss. This optimization implicitly follows a maximum likelihood estimation (MLE) framework under the assumption that the training data consists of independent and identically distributed samples drawn from the joint distribution of images and clean labels. However, when labels are noisy, the assumption of identically distributed assumption no longer holds, as the observed labels systematically deviate from clean labels, leading to performance degradation in the trained model.
@@
To address this issue, we maintain the MLE framework but reformulate it using a more suitable probabilistic model. Specifically, we introduce a model that explicitly accounts for the presence of noisy labels, which differ from the clean labels due to labeling errors. In practice, annotation errors tend to exhibit strong spatial correlations - mislabeling often occurs in contiguous regions. Variations in annotation criteria among experts, as well as changes in the underlying scene — such as the construction or demolition of buildings or alterations in vegetation — further reinforce this spatial dependency. Given this, we assume that label errors are not independent but instead exhibit spatial correlations among pixels.
@@
However, directly modeling spatial correlations poses significant computational challenges. To address this, we introduce a class of discrete distributions called the ELBO Computable Discrete Distribution, which enables the efficient optimization of the Evidence Lower Bound (ELBO). This discrete distribution is represented through continuous variables that follow a Gaussian distribution, where spatial correlations between pixels are expressed via a covariance matrix. Representing discrete variables through a Gaussian distribution successfully circumvents the intractability of summing over all possible realizations of the discrete variables. While the covariance matrix, whose number of elements scales quadratically with the number of pixels, introduces additional computational challenges in evaluating the ELBO, particularly in computing second-order statistics, its inverse, and its determinant. To overcome these computational intractabilities, we leverage the Kac-Murdock-Szegö (KMS) matrix [5, 17], which enables efficient computations necessary for ELBO evaluation. To validate the effectiveness of our approach, we conduct extensive empirical evaluations on multiple segmentation tasks. Our experimental results demonstrate that our method significantly improves robustness against label noise, particularly in scenarios with moderate to high levels of spatially correlated label noise. In summary, our contributions are as follows:
@@
    A probabilistic model with the ELBO Computable Discrete Distribution for learning with spatially correlated label noise:
    We propose a probabilistic model that explicitly accounts for noisy labels and their spatial correlations, which are prevalent in real-world applications, particularly in medical imaging and remote sensing. To address the computational challenges of modeling spatially correlated label noise, we introduce the ELBO Computable Discrete Distribution, which represents discrete variables via continuous variables following a Gaussian distribution. This approach circumvents the computational difficulties caused by correlated discrete random variables while preserving spatial dependencies in label errors.
@@
    Computing the covariance matrix in our probabilistic model is computationally intractable due to the number of elements scaling quadratically with the number of pixels. We overcome this challenge by leveraging the KMS matrix, which enables efficient computation of second-order statistics, matrix inversion, and determinant evaluation - key operations necessary for ELBO optimization.
@@ 2 Related Work
Learning with noisy labels has been a significant focus in the field of machine learning, particularly classification tasks [25, 40, 10, 34]. One research direction models the relationship between the clean label $y^{*}$ and the noisy label $y$ through a transition matrix [25, 40] that characterizes the probabilities from the clean label to the noisy label. For example, the transition probability from a clean label of $k$ to a noisy label of $j$ is represented as $p(y=j|y^{*}=k)$. Additionally, methods have been proposed to estimate an instance-dependent transition matrix [36, 2], where the transition probabilities depend on the instance $x$ also, denoted as $p(y=j|y^{*}=k,x)$. These classification methods can be adapted for semantic segmentation tasks; however, they treat label errors as being independent to each pixel, thereby ignoring the spatial correlations in label errors that may be present in annotations. This limitation is particularly critical in applications such as medical imaging [41, 29, 43] and remote sensing [24, 31, 21, 11], where different types of noise may occur.
Other lines of research includes work by Zhang et al. [44], which proposed a Tri-Network framework that trains three networks simultaneously, with each pair selecting reliable pixels based on their loss maps, thereby achieving robust learning even with coarse annotations. Liu et al. [22] introduced a mechanism to detect the moment when a network begins to overfit noisy labels. Gonzalez et al. [9] designed T-Loss based on a Student-t distribution to apply a logarithmic penalty on large errors, reducing undue influence from outlier pixels. Nonetheless, these methods do not take into account the spatial correlations in label errors that may occur within annotations.
Unlike the aforementioned related works, there are few studies that consider spatial correlations. Li et al. [20] leveraged superpixel segmentation to group pixels and smooth the network outputs within each superpixel, thereby indirectly incorporating spatial context. This method, however, relies heavily on the quality of the generated superpixels. More recently, Yao et al. [39] proposed an approach that explicitly models spatial correlations by applying a Markov process that consists of expansion and shrinkage steps of the annotation masks. Although this method achieves better performance, it is mainly limited to addressing noise along the boundaries. In real-world scenarios, segmentation labels are often corrupted by a diverse range of noise types, including omission noise, misalignment noise, boundary uncertainties, and other systematic labeling errors [15, 33, 21, 24, 14].
@@
3 Proposed Method
Figure 1: (a) Graphical model of the proposed model. Our proposed method introduces a latent Gaussian variable $\bm{\eta}$ that encodes spatial correlations of label error among adjacent pixels, thereby enabling a more realistic modeling of the dependency between pixels.
@@ 3 Proposed Method
3.1 Overview
In supervised learning, training a segmentation model typically involves optimizing the cross-entropy loss. The optimization of cross-entropy loss implicitly assumes that the labels in the training data are independent and identically distributed samples drawn from the joint distribution of images and clean labels. However, this assumption about the probabilistic model does not accurately reflect reality. To address this, we explicitly model the relationship between noisy labels and clean labels by introducing clean labels that are distinct from the given noisy labels within a probabilistic framework that captures their plausible dependencies. By learning the segmentation model parameters through marginal likelihood maximization, which integrates out unknown latent label errors, we aim to mitigate the impact of label noise. Since direct maximization of the marginal likelihood is intractable, we approximate it via ELBO maximization [3, 16], simultaneously estimating the approximate posterior distribution of label errors. By explicitly formulating a probabilistic model that accounts for label errors, the derived ELBO naturally leads to a well-justified loss function. Specifically, it includes the optimization of a cross-entropy loss computed using the estimated clean labels, where the contribution of each estimated clean label is weighted according to its associated uncertainty. The following describes the probabilistic model we have constructed.
Let $\mathcal{Y}$ be a set of labels, $\mathbf{x}\in\mathbb{R}^{HW\times C}$ be an input image, $\mathbf{y}\in\mathcal{Y}^{HW}$ be an observed label, $\mathbf{y}^{*}\in\mathcal{Y}^{HW}$ be a clean label, and $\bm{\theta}$ be the parameters of the semantic segmentation model. The likelihood of the observed label $p(\mathbf{y}|\mathbf{x},\bm{\theta})$ is defined as follows:
$p(\mathbf{y}|\mathbf{x},\bm{\theta})=\sum_{\mathbf{y}^{*}}p(\mathbf{y}|\mathbf{y}^{*})p(\mathbf{y}^{*}|\mathbf{x},\bm{\theta})=\sum_{\mathbf{y}^{*}}p(\mathbf{y}|\mathbf{y}^{*})\prod_{i}p_{\bm{\theta}}(y^{*}_{i}|\mathbf{x}).$ (1)
As mentioned in Section 1, real-world label errors such as omission and variation in expert annotation styles tend to exhibit spatial correlations. However, directly incorporating spatial correlations into discrete label variables makes computation intractable. To address this, we introduce a continuous variable, $\bm{\eta}\in\mathbb{R}^{HW}$, which is the logit of label error probability for each pixel that captures spatial correlations and indirectly imposes spatial dependencies on the labels as follows:
[…]
Figure 1(a) depicts a graphical model illustrating this assumption. Here, $p(\bm{\eta})$ is modeled as a Gaussian distribution $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})$, with its covariance matrix structured using a Kac-Murdock-Szegö (KMS) matrix [5, 17]. As we will see shortly, the combination of conditional independence $p(\mathbf{y}|\mathbf{y}^{*},\bm{\eta})=\prod_{i}p(y_{i}|y^{*}_{i},\eta_{i})$ and the KMS-structured covariance enables efficient optimization of the ELBO (Evidence Lower Bound) even though the distribution (2) itself cannot be directly evaluated due to the intractability of the integral. This approach ensures that even for high-dimensional problems with image-like structures of size $HW$, the computational complexity remains proportional to $HW$, making large-scale optimization feasible. We refer to the probabilistic model described in Eq.(2) as the ELBO computable discrete distribution.
By introducing a variational distribution $q(\mathbf{y}^{*},\bm{\eta}|\mathbf{y})=\left\{\prod_{i}q(y^{*}_{i}|y_{i},\eta_{i})\right\}q(\bm{\eta})$ as an approximate posterior and leveraging the ELBO computable discrete distribution, we can derive the ELBO as follows (see A.2 in Supplementary Material for details):

### v3

@@ Abstract
In semantic segmentation, the accuracy of models heavily depends on the high-quality annotations. However, in many practical scenarios, such as medical imaging and remote sensing, obtaining true annotations is not straightforward and usually requires significant human labor. Relying on human labor often introduces annotation errors, including mislabeling, omissions, and inconsistency between annotators. In the case of remote sensing, differences in procurement time can lead to misaligned ground-truth annotations. These label errors are not independently distributed, and instead usually appear in spatially connected regions where adjacent pixels are more likely to share the same errors. To address these issues, we propose an approximate Bayesian estimation based on a probabilistic model that assumes training data include label errors, incorporating the tendency for these errors to occur with spatial correlations between adjacent pixels. However, Bayesian inference for such spatially correlated discrete variables is notoriously intractable. To overcome this fundamental challenge, we introduce a novel class of probabilistic models, which we term the ELBO-Computable Correlated Discrete Distribution (ECCD). By representing the discrete dependencies through a continuous latent Gaussian field with a Kac-Murdock-Szegö (KMS) structured covariance, our framework enables scalable and efficient variational inference for problems previously considered computationally prohibitive. Through experiments on multiple segmentation tasks, we confirm that leveraging the spatial correlation of label errors significantly improves performance. Notably, in specific tasks such as lung segmentation, the proposed method achieves performance comparable to training with clean labels under moderate noise levels. Code is available at https://github.com/pfnet-research/Bayesian_SpatialCorr.
@@ 1 Introduction
Semantic segmentation, which involves classifying each pixel in an image into one of several classes, is a crucial task in computer vision. In supervised learning, the accuracy of segmentation models critically depends on the quality of the annotations in the training data. However, obtaining truly accurate pixel-level annotations is challenging in many practical applications. Even when expert annotators are employed, errors, omissions, and subjectivity in interpretation are inevitable, leading to inconsistencies in the datasets. In particular, high inter- and intra-annotator variability is widely reported in medical imaging, where experts may have differing interpretations of the same structures. For instance, several studies Zhang et al. (2020a); Lampert et al. (2016); Rädsch et al. (2023); Yang et al. (2023) highlight significant discrepancies among expert annotators; some delineate structures more generously, while others prefer more conservative annotations. Observer-dependent annotations exacerbate label noise in supervised learning. Label noise is also a critical issue in remote sensing, where determining ground truth labels often requires field surveys over large and sometimes inaccessible regions Frénay & Verleysen (2013); Pelletier et al. (2017); Foody (2002). Due to the logistical and economic challenges of large-scale ground truth collection, researchers frequently rely on automatic labeling systems, which may introduce systematic errors. Additionally, high-quality annotated datasets remain a critical bottleneck for supervised learning, particularly in remote sensing applications where annotations are often repurposed across different types of satellite images. For example, the OpenEarthMap dataset Xia et al. (2023) was created by manually annotating high-resolution optical satellite images for semantic segmentation. However, these annotations are sometimes reused for synthetic aperture radar (SAR) imagery, despite differences in resolution and capture conditions Huang et al. (2021); Zhang et al. (2021); Liu et al. (2023). Additionally, changes in artificial structures or variations in land cover further contribute to label inconsistencies Fritz et al. (2009); Huang et al. (2021).
@@
Various approaches have been proposed to mitigate the adverse effects of noisy labels. Some methods attempt to stop training early to prevent the network from overfitting to noise and generating unreliable pseudo-labels Liu et al. (2024; 2022), while others modify the loss function to be more robust against large errors Gonzalez-Jimenez et al. (2025). Although these techniques can reduce the influence of noisy annotations, they do not fundamentally address the core reason why the standard supervised learning framework fails in the presence of noisy labels.
@@
We propose a method that directly tackles the root cause of this issue. To understand this root cause, we first revisit the probabilistic assumptions underlying the standard cross-entropy loss. In supervised learning for segmentation models, training typically reduces to optimizing this loss function. This optimization implicitly follows a maximum likelihood estimation (MLE) framework under the assumption that the training data consists of independent and identically distributed samples drawn from the joint distribution of images and clean labels. However, when labels are noisy, the assumption of identically distributed assumption no longer holds, as the observed labels systematically deviate from clean labels, leading to performance degradation in the trained model.
@@
To address this issue, we maintain the MLE framework but generalize the underlying probabilistic model to make it more suitable for real-world scenarios with noisy labels. Specifically, we introduce a model that explicitly accounts for the presence of noisy labels, which differ from the clean labels due to labeling errors. In practice, annotation errors tend to exhibit strong spatial correlations - mislabeling often occurs in contiguous regions. Variations in annotation criteria among experts, as well as changes in the underlying scene — such as the construction or demolition of buildings or alterations in vegetation — further reinforce this spatial dependency. Given this, we assume that label errors are not independent but instead exhibit spatial correlations among pixels.
@@
However, directly modeling spatial correlations among discrete labels leads to an intractable marginal likelihood, a long-standing challenge in probabilistic modeling Koller & Friedman (2009); Sutton & McCallum (2012). Our core contribution is to resolve this challenge by proposing a novel class of tractable models, the ELBO-Computable Correlated Discrete Distribution (ECCD). This discrete distribution is represented through continuous variables that follow a Gaussian distribution, where spatial correlations between pixels are expressed via a covariance matrix. Representing discrete variables through a Gaussian distribution successfully circumvents the intractability of summing over all possible realizations of the discrete variables. While the covariance matrix, whose number of elements scales quadratically with the number of pixels, introduces additional computational challenges in evaluating the ELBO, particularly in computing second-order statistics, its inverse, and its determinant. To overcome these computational intractabilities, we leverage the Kac-Murdock-Szegö (KMS) matrix Fikioris (2018); Kac et al. (1953), which enables efficient computations necessary for ELBO evaluation. To validate the effectiveness of our approach, we conduct extensive empirical evaluations on multiple segmentation tasks. Our experimental results demonstrate that our method significantly improves robustness against label noise, particularly in scenarios with moderate to high levels of spatially correlated label noise. In summary, our contributions are as follows:
@@
    A Novel Class of Tractable Distributions for Correlated Discrete Variables:
    We introduce the ECCD, a new class of probabilistic models for tractable inference on high-dimensional discrete variables with stationary, exponentially decaying correlations—a structure common to spatial grids and temporal sequences. In this work, we apply the ECCD to learn from spatially correlated label noise in medical and remote sensing, where the model circumvents computational intractability by representing discrete dependencies via a continuous latent Gaussian field.
@@ A Novel Class of Tractable Distributions for Correlated Discrete Variables:
    The ECCD’s covariance matrix introduces computational challenges that scale quadratically with the number of pixels. We overcome this by leveraging the Kac-Murdock-Szegö (KMS) matrix. To the best of our knowledge, this is the first work to introduce the KMS matrix as a scalable computational tool for variational inference with correlated latent variables, enabling efficient computation of the key operations necessary for optimizing the ELBO of the ECCD.
@@ 2 Related Work
Learning with noisy labels has been a significant focus in the field of machine learning, particularly classification tasks Patrini et al. (2017); Yao et al. (2020); Han et al. (2018); Wei et al. (2020). One research direction models the relationship between the clean label $y^{*}$ and the noisy label $y$ through a transition matrix Patrini et al. (2017); Yao et al. (2020) that characterizes the probabilities from the clean label to the noisy label. For example, the transition probability from a clean label of $k$ to a noisy label of $j$ is represented as $p(y=j|y^{*}=k)$. Additionally, methods have been proposed to estimate an instance-dependent transition matrix Xia et al. (2020); Berthon et al. (2021), where the transition probabilities depend on the instance $x$ also, denoted as $p(y=j|y^{*}=k,x)$. These classification methods can be adapted for semantic segmentation tasks; however, they treat label errors as being independent to each pixel, thereby ignoring the spatial correlations in label errors that may be present in annotations. This limitation is particularly critical in applications such as medical imaging Zhang et al. (2020a); Shi & Wu (2021); Zhang et al. (2020b) and remote sensing Mnih & Hinton (2012); Song et al. (2022); Liu et al. (2024); Henry et al. (2021), where different types of noise may occur.
Other lines of research includes work by Zhang et al. Zhang et al. (2020c), which proposed a Tri-Network framework that trains three networks simultaneously, with each pair selecting reliable pixels based on their loss maps, thereby achieving robust learning even with coarse annotations. Liu et al. Liu et al. (2022) introduced a mechanism to detect the moment when a network begins to overfit noisy labels. Gonzalez et al. Gonzalez-Jimenez et al. (2025) designed T-Loss based on a Student-t distribution to apply a logarithmic penalty on large errors, reducing undue influence from outlier pixels. Nonetheless, these methods do not take into account the spatial correlations in label errors that may occur within annotations.
Unlike the aforementioned related works, there are few studies that consider spatial correlations. Li et al. Li et al. (2021) leveraged superpixel segmentation to group pixels and smooth the network outputs within each superpixel, thereby indirectly incorporating spatial context. This method, however, relies heavily on the quality of the generated superpixels. More recently, Yao et al. Yao et al. (2023) proposed an approach that explicitly models spatial correlations by applying a Markov process that consists of expansion and shrinkage steps of the annotation masks. Although this method achieves better performance, it is mainly limited to addressing noise along the boundaries. In real-world scenarios, segmentation labels are often corrupted by a diverse range of noise types, including omission noise, misalignment noise, boundary uncertainties, and other systematic labeling errors Jiménez & Decaestecker (2024); Vădineanu et al. (2022); Liu et al. (2024); Mnih & Hinton (2012); Jiang et al. (2021).
@@
M
3 Problem Formulation
Figure 1: (a) Graphical model of our proposed method, the ECCD. The model introduces a latent Gaussian variable $\bm{\eta}$ that encodes spatial correlations of label error among adjacent pixels, thereby enabling a more realistic modeling of the dependency between pixels.
@@ 3 Problem Formulation
3.1 Probabilistic Foundations of Standard Supervised Segmentation
In standard supervised learning for semantic segmentation, the training process is typically framed as the minimization of the cross-entropy loss. This objective function is not arbitrary; it is derived from the principle of Maximum Likelihood Estimation (MLE) under a specific, and often implicit, probabilistic model. This underlying probabilistic model makes two fundamental assumptions: (1) the provided annotations $\mathbf{y}$ are the true, noise-free clean labels $\mathbf{y}$$^{*}$, and (2) the label of each pixel $y_{i}$ is conditionally independent of the labels of other pixels given the input image $\mathbf{x}$.
Under these assumptions, the goal is to find the model parameters $\theta$ that maximizes the log-likelihood of the observed data:
$\log p_{\theta}(\mathbf{y}|\mathbf{x})=\log\prod_{i=1}^{HW}p_{\theta}(y_{i}|\mathbf{x})=\sum_{i=1}^{HW}\log p_{\theta}(y_{i}|\mathbf{x}),$ (1)
where $HW$ is the total number of pixels. The segmentation model $p_{\theta}(y_{i}|\mathbf{x})$ outputs a probability distribution over the set of possible classes $\mathcal{Y}$ for each pixel $i$. Let this distribution be represented by a vector of probabilities $\{p_{\theta}(y_{ik}|\mathbf{x})\}^{|\mathcal{Y}|}_{k=1}$. If we represent the ground-truth label $y_{i}$ as a one-hot vector, where $y_{ik}=1$ if pixel $i$ belongs to class $k$ and $y_{ik}=0$ otherwise, the log-likelihood for a single pixel can be written as $\sum_{k=1}^{|\mathcal{Y}|}y_{ik}\log p_{\theta}(y_{ik}|\mathbf{x})$. The total log-likelihood for the image is therefore:
$\mathcal{L}_{\text{MLE}}(\theta)=\sum_{i=1}^{HW}\sum_{k=1}^{|\mathcal{Y}|}y_{ik}\log p_{\theta}(y_{ik}|\mathbf{x}).$ (2)
Minimizing the negative of this quantity, $\mathcal{L}_{\text{MLE}}(\theta)$, is precisely equivalent to minimizing the standard multi-class cross-entropy loss. This derivation highlights that the validity of the cross-entropy loss is contingent on the assumption that the training labels are clean and pixel-wise independent samples from the true data distribution.
3.2 A General Probabilistic Model for Learning with Noisy Labels
In many practical scenarios, particularly in medical imaging and remote sensing, the assumption that observed labels $\mathbf{y}$ are identical to the true clean labels $\mathbf{y}^{*}$ does not hold. Annotations are often corrupted by noise, which can be spatially correlated. To address this, we generalize the probabilistic model by explicitly distinguishing between the observed labels $\mathbf{y}$ and the unobserved (latent) clean labels $\mathbf{y}^{*}$. This leads to the objective of maximizing the marginal log-likelihood, which integrates out the uncertainty over the unknown clean labels $\mathbf{y}^{*}$:
$\log p_{\theta}(\mathbf{y}|\mathbf{x})=\log\left(\sum_{\mathbf{y}^{*}}p_{\theta}(\mathbf{y},\mathbf{y}^{*}|\mathbf{x})\right)=\log\left(\sum_{\mathbf{y}^{*}}p(\mathbf{y}|\mathbf{y}^{*})p_{\theta}(\mathbf{y}^{*}|\mathbf{x})\right).$ (3)
Here, $p_{\theta}(\mathbf{y}^{*}|\mathbf{x})$ is the probability of the clean labels $\mathbf{y}^{*}$ given the image $\mathbf{x}$, which is modeled by the segmentation network, and $p(\mathbf{y}|\mathbf{y}^{*})$ represents the label noise process.
This formulation is a direct generalization of the standard MLE objective described in Section 3.1. To see this, consider the noise-free case where the noise model $p(\mathbf{y}|\mathbf{y}^{*})$ collapses to a Kronecker delta function $\delta(\mathbf{y}^{*},\mathbf{y})$, which equals 1 if and only if $\mathbf{y}^{*}=\mathbf{y}$, and is zero otherwise. In this scenario, the summation in Eq. (3) contains only one non-zero term, and the objective reduces precisely to maximizing $\log p_{\theta}(\mathbf{y}|\mathbf{x})$, which is the basis for the standard cross-entropy loss as shown in Eq. (1). However, the summation over all possible configurations of $\mathbf{y}^{*}$ is computationally intractable, as the number of terms $(|\mathcal{Y}|^{HW})$ grows exponentially with the number of pixels. One way to overcome this is to assume a pixel-independent noise process, $p(\mathbf{y}|\mathbf{y}^{*})=\prod_{i}p(y_{i}|y_{i}^{*})$, but this fails to capture real-world spatially correlated label noise, as mentioned in Section 1.
4 The Proposed ECCD Framework
4.1 A Tractable Variational Objective via the ECCD
To overcome the intractability of the marginal likelihood while modeling spatial correlations, we resort to variational inference and maximize its Evidence Lower Bound (ELBO). We introduce a continuous latent variable, $\bm{\eta}\in\mathbb{R}^{HW}$, which represents the logit of the label error probability for each pixel. This variable captures spatial correlations and indirectly imposes spatial dependencies on the labels as follows:
$p(\mathbf{y}|\mathbf{y}^{*})=\int p(\mathbf{y}|\mathbf{y}^{*},\bm{\eta})p(\bm{\eta})d\bm{\eta}=\int\left\{\prod_{i}p(y_{i}|y^{*}_{i},\eta_{i})\right\}p(\bm{\eta})d\bm{\eta}.$ (4)
Figure 1(a) depicts the graphical model illustrating this assumption. To fully specify this generative model, which is the central component of the ECCD, we define its constituent distributions. The prior $p(\bm{\eta})$ is modeled as a Gaussian, $p(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{\mu},\bm{\Sigma})$, with its covariance matrix structured using a Kac-Murdock-Szegö (KMS) matrix Fikioris (2018); Kac et al. (1953), which we detail in Section 4.2. The conditional noise model $p(y_{i}|y^{*}_{i},\eta_{i})$ is defined as:
$p(y_{ik}=1|y_{ic}^{*}=1,\eta_{i})=\delta_{kc}(1-r(\eta_{i}))+(1-\delta_{kc})r(\eta_{i})W_{kc},$ (5)
where $\delta_{kc}$ is the Kronecker delta, $r(\cdot)$ is the sigmoid function that maps the latent variable $\eta_{i}$ to a label error probability, and $W$ is a learnable class transition matrix that specifies the probability distribution over incorrect labels when an error occurs, with its elements satisfying $W\in[0,1]^{{|\mathcal{Y}|}\times{|\mathcal{Y}|}},\ W_{cc}=0$ and $\sum_{k=1}^{|\mathcal{Y}|}W_{kc}=1$, for each true class $c\in\mathcal{Y}$.
To derive a tractable objective for this model, we introduce a variational distribution $q(\mathbf{y}^{*},\bm{\eta}|\mathbf{y})$ to approximate the true posterior. We assume a factorized form symmetric to the generative model: $q(\mathbf{y}^{*},\bm{\eta}|\mathbf{y})=q(\bm{\eta})\prod_{i}q(y^{*}_{i}|y_{i},\eta_{i})$. The variational posterior for $\bm{\eta}$ is also a Gaussian, $q(\bm{\eta})=\mathcal{N}(\bm{\eta}|\bm{m},\bm{\Gamma})$, and the conditional posterior for the clean labels is parameterized as:
$q(y_{ik}^{*}=1|y_{ic}=1,\eta_{i})=\delta_{kc}(1-r(\eta_{i}))+(1-\delta_{kc})r(\eta_{i})V_{kc},$ (6)
where $V\in[0,1]^{{|\mathcal{Y}|}\times{|\mathcal{Y}|}}$ is another learnable class transition matrix modeling the inverse noise process (i.e., inferring the clean label from the noisy one), with analogous constraints ($V_{cc}=0$ and $\sum_{k=1}^{|\mathcal{Y}|}V_{kc}=1$ for each noisy class $c\in\mathcal{Y}$). With the generative model and variational posterior fully defined, we can derive the Evidence Lower Bound (ELBO) as follows (see Appendix A.2 for details):